### v1

# Discord-type quantum correlations in axially symmetric spin-(1/2, $S$) systems

Abstract
A mixed spin-$(1/2,S)$ system with arbitrary spin $S$ and interactions satisfying the U(1) axial symmetry is considered. Compact explicit forms of the local quantum uncertainty (LQU) and local quantum Fisher information (LQFI) are obtained directly through the entries and eigenvalues of the density matrix. Using these forms, we perform a comparative study of discord-type quantum correlations LQU and LQFI for the system at thermal equilibrium. An increase in quantum correlations with increasing spin length $S$ is discovered. Moreover, we find that as the system cools, quantum correlations can undergo one or more abrupt transitions while the temperature changes smoothly.
Quantum Fisher information, Wigner-Yanase skew information, Qubit-qudit system, Axial U(1) symmetry, Sudden transitions
pacs:
03.65.Aa, 03.65.Ud, 03.67.-a, 75.10.Jm
I Introduction
Quantum information technologies are rapidly entering all aspects of human life. The current state of quantum computing is called the Noisy Intermediate-Scale Quantum (NISQ) era [1, 2, 3]. Today’s quantum computers are limited to hundreds of qubits and have relatively short coherence times and significant errors [2, 4]. In general, they are far from universal, fault-tolerant quantum computers, which are believed to make it possible to effectively simulate any processes occurring in Nature.
Nevertheless, existing quantum computers can already solve important real-world problems. In search of the quantum advantage of these devices, NISQ algorithms have been proposed for applications in various disciplines, including physics, quantum chemistry, materials science, machine learning, etc. The goal of such algorithms is to exploit the limited available resources to perform classically complex tasks.
At the heart of quantum computing lie extraordinary correlations between subsystems of a quantum-mechanical system. Until the 21st century, the quantum entanglement was considered to be the only ingredient of quantum properties leading to such correlations. It manifests itself in the Einstein-Podolsky-Rosen gedanken (thought) experiment, Bell’s inequality test, quantum cryptography, superdense coding, teleportation, and so on [5, 6] (see also review articles [7, 8, 9, 10]).
Quantum entanglement was quantified in 1996, first for pure states [11, 12], and then for mixed states [13]. According to the accepted definition, the entanglement of a bipartite pure state is the von Neumann entropy of either of the two subsystems. The entanglement (of formation) of a bipartite mixed state is defined as the minimum entanglement of an ensemble over all ensembles realizing the mixed state; using this definition, Wootters derived an exact formula for general two-qubit systems [14]. Another important measure of the quantum entanglement was proposed in 2002 by Vidal and Werner [15]. This measure, called negativity, originated from the Peres-Horodecki PPT-criterion [16, 17]. Moreover, many measures have been introduced so far to quantify the entanglement of bipartite and multipartite systems, each of which is efficient for specific quantum states [18, 19].
At one time it was believed that quantum entanglement was the only resource of quantum speedup in quantum computation. However, in 1998, Knill and Laflamme showed, using the model of deterministic quantum computation with one pure qubit (DQC1) [20], that the computation can achieve an exponential improvement in efficiency over classical computers even without containing much entanglement. In 2000-2001, $\rm\dot{Z}$urek et al. developed the concept of quantum discord – “a measure of the quantumness of correlations” [21, 22]. Simultaneously and independently, Vedral et al. [23, 24] proposed a measure for the purely classical correlation, which, after subtracting it from the total correlation, led to the same amount of quantum correlation as the discord. Then, Datta et al. [25, 26] calculated discord in the Knill-Laflamme DQC1 model and proved that it scales with the quantum efficiency, while entanglement remains vanishingly small throughout the computation. This attracted a lot of attention to the new measure of quantum correlation [27, 28, 29, 30, 31, 32].
Quantum discord and entanglement are the same for the pure quantum states. However, discord can exist in separable mixed states, i.e., when quantum entanglement is identically equal to zero. The set of separable states possesses a nonzero volume in the whole Hilbert space of a system [33] (it is a necessary condition for the arising of entanglement sudden death (ESD) effect [34]), whereas the set of states with zero discord, vise versa, is negligibly small [35]. This circumstance alone sharply distinguishes discord from entanglement. Moreover, numerous theoretical and experimental investigations of different quantum systems have clearly shown that while the quantum entanglement and discord measure the same thing – the quantum correlation – but as a matter of fact, discrepancies in quantitative and even qualitative behavior are large [36, 37, 38]. Discord and entanglement behave differently even for the simplest mixed states, namely the Werner and Bell-diagonal ones (see, e.g., [39]). This has led many to talk about entanglement and discord as different types of quantum correlations.
The subsequent proposals with more and more new measures of quantum correlations [40, 41] caused a dilemma: should each measure be attributed to its own correlation, or should it be argued that there is only one quantum correlation, but the methods of describing it are not adequate? The physicists community now prefers to talk about entanglement and discord-like quantum correlations [42]. As is customary for brevity, we will also refer to the various measures of quantum correlation as “quantum correlations”. Nevertheless, the quantum correlation is one, but now there are only different measures for it, which are still imperfect. We would like to have such measures of quantum correlations that would be useful for estimating the values of the speedup of quantum computing, the efficiency of quantum heat engines, etc. Let the measures be different, but the results must be the same. As, for example, there are various formulations of quantum mechanics which differ dramatically in mathematical and conceptual overview, yet each one makes identical predictions for all experimental results [43]. The existing difficulties only increase interest in quantum correlations because “who knows, studying the nature of correlations in the world around us might even help us catch a glimpse of the theory that comes to supersede quantum physics” [29].
The concept of quantum discord is based on local measurements and optimization that were applied to the mutual information. This idea was then used for other physical or information quantities, and to date, a large number of discord-like measures of quantum correlation have been introduced, such as quantum deficit, geometric quantum discord, measurement induced disturbance and so on [29, 40, 41]. The local quantum uncertainty (LQU) and local quantum Fisher information (LQFI), based respectively on the Wigner-Yanase and quantum Fisher information, belong to this type of measures [44, 45].
[…]
Nonclassical correlations in various special cases of the mixed spin-$(1/2,S)$ model were considered in a number of works. For example, the quantum entanglement, quantified by negativity, was investigated in Refs. [51, 52, 53, 54, 55]. Recently, two studies have been carried out on quantum correlations in hybrid qubit-qutrit systems using negativity, as well as LQU and LQFI [56, 57]. In the paper [58], we provided exact formulas of LQU and LQFI for general qubit-qubit X states. Extending this approach to arbitrary spin $S$, we evaluate here the exact analytical expressions for LQU and LQFI in the case of $(1/2,S)$ axisymmetric states and examine the behavior of thermal quantum correlations for different choices of Hamiltonian parameters.
The present paper is structured as follows. In the next section, we establish the general form of the axially symmetric Hamiltonian and density matrix. Section III is devoted to the diagonalization of the density matrix. In Section IV, we derive formulas suitable for actually calculating the LQU and LQFI correlations. The spin-(1/2,$S$) system at thermal equilibrium is considered in Sect. V. Section VI includes the numerical results and discussion. Finally, concluding remarks close this paper.
II Hamiltonian and density matrix
[…]
${\cal S}_{z}=s_{z}\otimes I_{2S+1}+I_{2}\otimes S_{z}={\rm diag}\,[1/2+S,1/2+S-1,\ldots,1/2-S,-1/2+S,\ldots,-1/2-S+1,-1/2-S],$ (1)
where $I_{2}$ and $I_{2S+1}$ are the identity operators of the second and ($12S+$ 1)-th orders, respectively. The matrix (1) of $2(2S+1)$-th order is diagonal, and let its eigenvalues be numbered from 0 to $14S+$ 1. Except for the extreme eigenvalues ±$\pm(1/2+S)$, the remaining $4S$ internal eigenvalues of the matrix (1) are two-fold degenerate: the $k$-th and $(2S+k)$-th ones are equal to each other ($k=1,\ldots,2S$).
The most general matrix which commutes with the $z$-component of total spin (1) can be written as
${\cal A}=\left(\begin{array}[]{cccccccccc}a_{00}&&&&&&&&&\\ &a_{11}&0&\ldots&0&a_{1,2S+1}&0&\ldots&0&\\ &0&a_{22}&\ldots&0&0&a_{2,2S+2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&a_{2S,2S}&0&0&\ldots&a_{2S,4S}&\\ &a_{2S+1,1}&0&\ldots&0&a_{2S+1,2S+1}&0&\ldots&0&\\ &0&a_{2S+2,2}&\ldots&0&0&a_{2S+2,2S+2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&a_{4S,2S}&0&0&\ldots&a_{4S,4S}&\\ &&&&&&&&&a_{4S+1,4S+1}\end{array}\right).$ (2)
This matrix has a block-diagonal structure $(1\times 1)\oplus(4S\times 4S)\oplus(1\times 1)$, where the interior $4S\times 4S$ subblock is sparse, and its nonzero entries are located only on the main diagonal and two sub-diagonals. Remarkably, the set of matrices of such a type are algebraically closed: their sums and products preserve the same form. Below, for brevity, such matrices generated by the axial symmetry will be called $\cal A$ matrices.
If the matrix (2) is Hermitian, as in the case of Hamiltonians or density matrices, it may have $2(4S+1)$ real independent parameters. Expansion of the matrix ${\cal H}$ into the spin-1/2 and spin-$S$ matrices and their products and powers allows us to clarify the structure and physical meaning of interactions in the system (see below).
For a given Hamiltonian $\cal H$ in the $\cal A$ form, the density matrix $\rho$ can be obtained from solving some master equations. Let the initial state has the $\cal A$ structure and the dynamics of the system is governed, for example, by the quantum Liouville equation. Then, the $\cal A$ form of the density matrix will be preserved throughout the entire evolution process due to the above-mentioned property of the algebraic closedness of $\cal A$ matrices. The $\cal A$-form of the density matrix will also be realized for systems in thermal equilibrium, since $\rho$ and $\cal H$ are related by functional dependence. In these and similar cases, the density matrix under examination can be written as
$\rho=\left(\begin{array}[]{clccccccrc}p_{0}&&&&&&&&&\\ &a_{1}&0&\ldots&0&u_{1}&0&\ldots&0&\\ &0&a_{2}&\ldots&0&0&u_{2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&a_{2S}&0&0&\ldots&u_{2S}&\\ &u_{1}^{*}&0&\ldots&0&a_{2S+1}&0&\ldots&0&\\ &0&u_{2}^{*}&\ldots&0&0&a_{2S+2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&u_{2S}^{*}&0&0&\ldots&a_{4S}&\\ &&&&&&&&&p_{4S+1}\end{array}\right).$ (3)
Below, we restrict ourselves to these $\cal A$ quantum states.
III Diagonal representation of the density matrix
III.1 Diagonalization of the matrix $\rho$
Let us turn to the diagonalization of matrix (3). Since the density matrix $\rho$ commutes with the diagonal matrix ${\cal S}_{z}$, it has nonzero matrix elements only between states with the same eigenvalue of ${\cal S}_{z}$. Here, ${\cal S}_{z}$ plays a role of “constant or integral of motion”; in this connection see, e.g., Chapter 1 in two books [59, 60]. As mentioned above, ${\cal S}_{z}$ has two non-degenerate and $4S$ doubly degenerate eigenvalues. Therefore, the density matrix (3) actually consists of two one-by-one and $2S$ two-by-two subblocks.111 This will be obvious if we reorder the vectors of the original basis so that the $k$th and $(2S+k)$th vectors become neighbors, then the density matrix (3) takes an explicit quasi-diagonal form: $\hat{\rho}=[p_{0},\left(\begin{array}[]{cc}a_{1}&u_{1}\\ u_{1}^{*}&a_{2S+1}\\ \end{array}\right),\ldots,\left(\begin{array}[]{cc}a_{2S}&u_{2S}\\ u_{2S}^{*}&a_{4S}\\ \end{array}\right),p_{4S+1}].$ This form is implemented for the system $(S,1/2)$. The first and $(4S+1)$th “subblocks” have sizes $1\times 1$, i.e. they are ready eigenvalues. The remaining $2S$ subblocks are $2\times 2$-dimensional, and their diagonalization is easily accomplished by directly calculating the eigenvectors of the corresponding subblocks (see, for example, [61]). Eigenvectors are constructed using the $k$th and $(2S+k)$th pair of vectors of the original basis. Indeed, a suitable fragment extracted from the inner $4S\times 4S$ block of the density matrix (3) reads
$\left(\begin{array}[]{ccc}a_{k}&\ldots&u_{k}\\ \vdots&\ddots&\vdots\\ u_{k}^{*}&\ldots&a_{2S+k}\\ \end{array}\right),\qquad k=1,\ldots,2S$ (4)
and then the corresponding eigenvectors are given as
$|k\rangle=\left(\begin{array}[]{c}{\tilde{q}}_{k}\\ \vdots\\ {\tilde{u}}_{k}^{*}\\ \end{array}\right),\qquad|2S+k\rangle=\left(\begin{array}[]{c}{\tilde{u}}_{k}\\ \vdots\\ -{\tilde{q}}_{k}\\ \end{array}\right),\qquad$ (5)
where
${\tilde{q}_{k}}=q_{k}/\sqrt{q_{k}^{2}+|u_{k}|^{2}},\qquad{\tilde{u}_{k}}=u_{k}/\sqrt{q_{k}^{2}+|u_{k}|^{2}},\qquad q_{k}=\frac{1}{2}\Big{(}a_{k}-a_{2S+k}+\sqrt{(a_{k}-a_{2S+k})^{2}+4|u_{k}|^{2}}\Big{)}.$ (6)
This makes it possible to immediately write down the desired transformation, which diagonalizes the density matrix (3):
$R=\left(\begin{array}[]{clccccccrc}1&&&&&&&&&\\ &{\tilde{q}}_{1}&0&\ldots&0&{\tilde{u}}_{1}&0&\ldots&0&\\ &0&{\tilde{q}}_{2}&\ldots&0&0&{\tilde{u}}_{2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&{\tilde{q}}_{2S}&0&0&\ldots&{\tilde{u}}_{2S}&\\ &{\tilde{u}}_{1}^{*}&0&\ldots&0&-{\tilde{q}}_{1}&0&\ldots&0&\\ &0&{\tilde{u}}_{2}^{*}&\ldots&0&0&-{\tilde{q}}_{2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&{\tilde{u}}_{2S}^{*}&0&0&\ldots&-{\tilde{q}}_{2S}&\\ &&&&&&&&&1\end{array}\right),$ (7)
where ${\tilde{q}}_{k}$ and ${\tilde{u}}_{k}$ ($k=1,\ldots,2S$) are defined by Eq. (6). Note that $R$ is the $\cal A$ matrix. Moreover, it is clear that $R^{\dagger}=R$ and $R^{\dagger}R=I$, so this diagonalizing transformation is both Hermitian and unitary.
Using Eqs. (3) and (7) we find
$\rho^{\prime}\equiv R\rho R=[p_{0},p_{1},\ldots,p_{2S},p_{2S+1},\ldots,p_{4S},p_{4S+1}],$ (8)
where
$p_{k,2S+k}=\frac{1}{2}\Big{(}a_{k}+a_{2S+k}\pm\sqrt{(a_{k}-a_{2S+k})^{2}+4|u_{k}|^{2}}\Big{)}$ (9)
are the eigenvalues of the density matrix (3). From here and Eq. (6) it follows that
$q_{k}=p_{k}-a_{2S+k}$ (10)
and
${\tilde{q}}_{k}^{2}=\frac{1}{2}\Big{(}1+\frac{a_{k}-a_{2S+k}}{p_{k}-p_{2S+k}}\Big{)},\qquad{\tilde{u}}_{k}^{2}=\frac{1}{2}\Big{(}1-\frac{a_{k}-a_{2S+k}}{p_{k}-p_{2S+k}}\Big{)},$ (11)
where as before $k=1,\ldots,2S$.
III.2 Spin matrices $\sigma_{\mu}\otimes I_{2S+1}$ in $\rho^{\prime}$-representation
Let us transform the local Pauli matrices $\sigma_{\mu}\otimes I_{2S+1}$ ($\mu=x,y,z$) of the particle $A$ into a diagonal representation of the density operator. Using Eq. (7) we get
$R(\sigma_{x}\otimes I_{2S+1})R=\left(\begin{array}[]{cccccccccccc}&{\tilde{u}}_{1}^{*}&&&&&-{\tilde{q}}_{1}&&&&&\\ {\tilde{u}}_{1}&0&{\tilde{q}}_{1}{\tilde{u}}_{2}^{*}&\ldots&0&0&0&-{\tilde{q}}_{1}{\tilde{q}}_{2}&\ldots&0&0&\\ &{\tilde{q}}_{1}{\tilde{u}}_{2}&0&\ldots&0&0&{\tilde{u}}_{1}{\tilde{u}}_{2}&0&\ldots&0&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\\ &0&0&\ldots&0&{\tilde{q}}_{2S-1}{\tilde{u}}_{2S}^{*}&0&0&\ldots&0&-{\tilde{q}}_{2S-1}{\tilde{q}}_{2S}&\\ &0&0&\ldots&{\tilde{q}}_{2S-1}{\tilde{u}}_{2S}&0&0&0&\ldots&{\tilde{u}}_{2S-1}{\tilde{u}}_{2S}&0&{\tilde{q}}_{2S}\\ -{\tilde{q}}_{1}&0&{\tilde{u}}_{1}^{*}{\tilde{u}}_{2}^{*}&\ldots&0&0&0&-{\tilde{q}}_{2}{\tilde{u}}_{1}^{*}&\ldots&0&0&\\ &-{\tilde{q}}_{1}{\tilde{q}}_{2}&0&\ldots&0&0&-{\tilde{q}}_{2}{\tilde{u}}_{1}&0&\ldots&0&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\\ &0&0&\ldots&0&{\tilde{u}}_{2S-1}^{*}{\tilde{u}}_{2S}^{*}&0&0&\ldots&0&-{\tilde{q}}_{2S}{\tilde{u}}_{2S-1}^{*}&\\ &0&0&\ldots&-{\tilde{q}}_{2S-1}{\tilde{q}}_{2S}&0&0&0&\ldots&-{\tilde{q}}_{2S}{\tilde{u}}_{2S-1}&0&{\tilde{u}}_{2S}^{*}\\ &&&&&{\tilde{q}}_{2S}&&&&&{\tilde{u}}_{2S}&\end{array}\right),$ (12)
$R(\sigma_{y}\otimes I_{2S+1})R\!=\!\!\left(\begin{array}[]{cccccccccccc}&-i{\tilde{u}}_{1}^{*}&&&&&i{\tilde{q}}_{1}&&&&&\\ i{\tilde{u}}_{1}&0&-i{\tilde{q}}_{1}{\tilde{u}}_{2}^{*}&\ldots&0&0&0&i{\tilde{q}}_{1}{\tilde{q}}_{2}&\ldots&0&0&\\ &i{\tilde{q}}_{1}{\tilde{u}}_{2}&0&\ldots&0&0&i{\tilde{u}}_{1}{\tilde{u}}_{2}&0&\ldots&0&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\\ &0&0&\ldots&0&-i{\tilde{q}}_{2S-1}{\tilde{u}}_{2S}^{*}&0&0&\ldots&0&i{\tilde{q}}_{2S-1}{\tilde{q}}_{2S}&\\ &0&0&\ldots&i{\tilde{q}}_{2S-1}{\tilde{u}}_{2S}&0&0&0&\ldots&i{\tilde{u}}_{2S-1}{\tilde{u}}_{2S}&0&-i{\tilde{q}}_{2S}\\ -i{\tilde{q}}_{1}&0&-i{\tilde{u}}_{1}^{*}{\tilde{u}}_{2}^{*}&\ldots&0&0&0&i{\tilde{q}}_{2}{\tilde{u}}_{1}^{*}&\ldots&0&0&\\ &-i{\tilde{q}}_{1}{\tilde{q}}_{2}&0&\ldots&0&0&-i{\tilde{q}}_{2}{\tilde{u}}_{1}&0&\ldots&0&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\\ &0&0&\ldots&0&-i{\tilde{u}}_{2S-1}^{*}{\tilde{u}}_{2S}^{*}&0&0&\ldots&0&i{\tilde{q}}_{2S}{\tilde{u}}_{2S-1}^{*}&\\ &0&0&\ldots&-i{\tilde{q}}_{2S-1}{\tilde{q}}_{2S}&0&0&0&\ldots&-i{\tilde{q}}_{2S}{\tilde{u}}_{2S-1}&0&-i{\tilde{u}}_{2S}^{*}\\ &&&&&i{\tilde{q}}_{2S}&&&&&i{\tilde{u}}_{2S}&\end{array}\right)$ (13)
and
$R(\sigma_{z}\otimes I_{2S+1})R=\left(\begin{array}[]{clccccccrc}1&&&&&&&&&\\ &{\tilde{q}}_{1}^{2}-|{\tilde{u}}_{1}|^{2}&0&\ldots&0&2{\tilde{q}}_{1}{\tilde{u}}_{1}&0&\ldots&0&\\ &0&{\tilde{q}}_{2}^{2}-|{\tilde{u}}_{2}|^{2}&\ldots&0&0&2{\tilde{q}}_{2}{\tilde{u}}_{2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&{\tilde{q}}_{2S}^{2}-|{\tilde{u}}_{2S}|^{2}&0&0&\ldots&2{\tilde{q}}_{2S}{\tilde{u}}_{2S}&\\ &2{\tilde{q}}_{1}{\tilde{u}}_{1}^{*}&0&\ldots&0&-{\tilde{q}}_{1}^{2}+|{\tilde{u}}_{1}|^{2}&0&\ldots&0&\\ &0&2{\tilde{q}}_{2}{\tilde{u}}_{2}^{*}&\ldots&0&0&-{\tilde{q}}_{2}^{2}+|{\tilde{u}}_{2}|^{2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&2{\tilde{q}}_{2S}{\tilde{u}}_{2S}^{*}&0&0&\ldots&-{\tilde{q}}_{2S}^{2}+|{\tilde{u}}_{2S}|^{2}&\\ &&&&&&&&&-1\end{array}\right).$ (14)
Unlike the first two matrices, the latter has the $\cal A$ form. The entries $\langle m|\sigma_{\mu}\otimes I_{2S+1}|n\rangle$ of the matrices (12) – (14) will be needed in the next section.
IV Quantum correlations
IV.1 LQU
The total uncertainty of an observable $H$ in the quantum-mechanical state $\rho$ is usually expressed by the variance
${\rm Var}(\rho,H)=\langle H^{2}\rangle_{\rho}-\langle H\rangle_{\rho}^{2}.$ (15)
On the other hand, the quantum contribution to the total statistical error may be reliable quantify via the Wigner-Yanase skew information [62, 63, 64]
${\cal I}(\rho,H)=-\frac{1}{2}{\rm Tr}[\sqrt{\rho},H]^{2},$ (16)
where [.,.][.,.][ . , . ] denotes the commutator. Notice that the skew information (16) is not grater than the variance (15), namely
${\cal I}(\rho,H)\leq{\rm Var}(\rho,H),$ (17)
where the equality is achieved for pure states when classical ignorance does not occur [44]. This makes it possible to introduce a discord-type measure $\cal U$ (also called LQU) of quantum correlations in any bipartite system $AB$ as follows [44]
${\cal U}(\rho)=\min_{H_{A}}{\cal I}(\rho,H_{A}),$ (18)
in which the minimum is taken over all local observables $H_{A}$ on the subsystem $A$. It is worth mentioning that the LQU is a genuine quantifier of quantum correlations, and it has been shown that the LQU meets all the physical conditions of a measure of quantum correlations.
Importantly, the authors [44] were able to perform optimization for qubit-qudit systems and presented the measure (18) in the simple form
${\cal U}=1-\lambda_{max}^{(W)},$ (19)
where $\lambda_{max}^{(W)}$ denotes the maximum eigenvalue of the $3\times 3$ symmetric matrix $W$ whose entries are
$W_{\mu\nu}={\rm Tr}\{\rho^{1/2}(\sigma_{\mu}\otimes{\rm I}_{2S+1})\rho^{1/2}(\sigma_{\nu}\otimes{\rm I}_{2S+1})\}$ (20)
with $\mu,\nu=x,y,z$ and $\sigma_{x,y,z}$ are the Pauli matrices as before. This opens up the possibility of real calculating LQU.
Using Eqs. (12)–(14), we find that the matrix $W$ is diagonal, $W_{yy}=W_{xx}$ and
$\frac{1}{2}W_{xx}={\tilde{q}}_{1}^{2}\sqrt{p_{0}p_{2S+1}}+{\tilde{q}}_{2S}^{2}\sqrt{p_{2S}p_{4S+1}}+|{\tilde{u}}_{1}|^{2}\sqrt{p_{0}p_{1}}+|{\tilde{u}}_{2S}|^{2}\sqrt{p_{4S}p_{4S+1}}$
$+\sum_{k=1}^{2S-1}\big{[}{\tilde{q}}_{k}^{2}{\tilde{q}}_{k+1}^{2}\sqrt{p_{k}p_{2S+1+k}}+{\tilde{q}}_{k}^{2}|{\tilde{u}}_{k+1}|^{2}\sqrt{p_{k}p_{k+1}}+|{\tilde{u}}_{k}|^{2}{\tilde{q}}_{k+1}^{2}\sqrt{p_{2S+k}p_{2S+1+k}}+|{\tilde{u}}_{k}|^{2}|{\tilde{u}}_{k+1}|^{2}\sqrt{p_{k+1}p_{2S+k}}\big{]}$ (21)
and
$W_{zz}=p_{0}+p_{4S+1}+\sum_{k=1}^{2S}\big{[}(p_{k}+p_{2S+k})({\tilde{q}}_{k}^{2}-|{\tilde{u}}_{k}|^{2})^{2}+8{\tilde{q}}_{k}^{2}|{\tilde{u}}_{k}|^{2}\sqrt{p_{k}p_{2S+k}}\big{]}.$ (22)
Therefore, LQU would be given by
${\cal U}=\min{\{{\cal U}_{0},{\cal U}_{1}\}},$ (23)
where two branches (sub-functions) are ${\cal U}_{0}=1-W_{zz}$ and ${\cal U}_{1}=1-W_{xx}$.
Substituting expressions (11) into Eqs. (IV.1) and (22), we arrive at the final formulas for the branches
${\cal U}_{0}=\sum_{k=1}^{2S}\Bigg{[}(\sqrt{p_{k}}-\sqrt{p_{2S+k}})^{2}-\frac{(a_{k}-a_{2S+k})^{2}}{(\sqrt{p_{k}}+\sqrt{p_{2S+k}})^{2}}\Bigg{]}$ (24)
and
${\cal U}_{1}=1-2\Bigg{[}\frac{a_{2S+1}+\sqrt{p_{1}p_{2S+1}}}{\sqrt{p_{1}}+\sqrt{p_{2S+1}}}\sqrt{p_{0}}+\frac{a_{2S}+\sqrt{p_{2S}p_{4S}}}{\sqrt{p_{2S}}+\sqrt{p_{4S}}}\sqrt{p_{4S+1}}+\sum_{k=1}^{2S-1}\frac{(a_{k}+\sqrt{p_{k}p_{2S+k}})(a_{2S+1+k}+\sqrt{p_{k+1}p_{2S+1+k}})}{(\sqrt{p_{k}}+\sqrt{p_{2S+k}})(\sqrt{p_{k+1}}+\sqrt{p_{2S+1+k}})}\Bigg{]}.$ (25)
It is remarkable that both quantities ${\cal U}_{0}$ and ${\cal U}_{1}$ are expressed directly only through the matrix elements of the density matrix (3) and its eigenvalues, therefore, the same is true for the measure ${\cal U}$.
IV.2 LQFI
Quantum Fisher information (QFI) is the cornerstone of quantum estimation theory [65, 66, 67, 68, 69]. Suppose there is unitary evolution of the quantum state $\varrho=e^{iH\epsilon}\rho e^{-iH\epsilon}$ with some observable $H$, and the parameter $\epsilon$ wanted to be estimated. The QFI can be defined as follows
$F(\varrho,H)=\frac{1}{4}{\rm Tr}(\varrho L_{\epsilon}^{2}),$ (26)
where $L_{\epsilon}$ is the symmetric logarithmic derivative (SLD) operator satisfying the Lyapunov equation (known from control theory)
$\frac{\partial\varrho}{\partial\epsilon}=\frac{1}{2}(\varrho L_{\epsilon}+L_{\epsilon}\varrho).$ (27)
[…]
$\Delta\epsilon\geq\frac{1}{\sqrt{NF(\varrho,H)}},$ (28)
where $N$ is the number of measurements.
Another application of QFI is to use it as a quantifier of quantum correlations. Girolami et al. [45] showed that the quantity
${\cal F}(\rho)=\min_{H_{A}}F(\rho,H_{A}),$ (29)
where $H_{A}$ (as in the case of LQU) is a local observable, again satisfies all the necessary criteria to be defined as a measure of quantum correlations. Note that here the (local) QFI is minimized, unlike the case of the lower bound (28). The measure $\cal F$ is called interferometric power [45] or LQFI [70, 71]. Moreover, if the subsystem $A$ is a qubit, then minimization in (29) can be done and LQFI is given as [45]
${\cal F}=1-\lambda_{max}^{(M)},$ (30)
where $\lambda_{max}^{(M)}$ is the largest eigenvalue of the real symmetric $3\times 3$ matrix $M$ with entries [72]
$M_{\mu\nu}=\sum_{{m,n};\ {p_{m}+p_{n}\neq 0}}\frac{2p_{m}p_{n}}{p_{m}+p_{n}}\langle m|\sigma_{\mu}\otimes{\rm I}_{2S+1}|n\rangle\langle n|\sigma_{\nu}\otimes{\rm I}_{2S+1}|m\rangle.$ (31)
From a programming point of view, the right side of this equation represents two nested loops.
Again using Eqs. (12)–(14), we find that the matrix $M$ is diagonal, $M_{yy}=M_{xx}$, and
$\frac{1}{4}M_{xx}$ $=$ $\frac{p_{0}p_{2S+1}}{p_{0}+p_{2S+1}}{\tilde{q}}_{1}^{2}+\frac{p_{2S}p_{4S+1}}{p_{2S}+p_{4S+1}}{\tilde{q}}_{2S}^{2}+\frac{p_{0}p_{1}}{p_{0}+p_{1}}|{\tilde{u}}_{1}|^{2}+\frac{p_{4S}p_{4S+1}}{p_{4S}+p_{4S+1}}|{\tilde{u}}_{2S}|^{2}+\sum_{k=1}^{2S-1}\Big{[}\frac{p_{k}p_{2S+1+k}}{p_{k}+p_{2S+1+k}}{\tilde{q}}_{k}^{2}{\tilde{q}}_{k+1}^{2}$ (32)
$+$ $\frac{p_{k}p_{k+1}}{p_{k}+p_{k+1}}{\tilde{q}}_{k}^{2}|{\tilde{u}}_{k+1}|^{2}+\frac{p_{2S+k}p_{2S+1+k}}{p_{2S+k}+p_{2S+1+k}}|{\tilde{u}}_{k}|^{2}{\tilde{q}}_{k+1}^{2}+\frac{p_{k+1}p_{2S+k}}{p_{k+1}+p_{2S+k}}|{\tilde{u}}_{k}|^{2}|{\tilde{u}}_{k+1}|^{2}\Big{]}$
and
$M_{zz}=p_{0}+p_{4S+1}+\sum_{k=1}^{2S}\big{[}(p_{k}+p_{2S+k})({\tilde{q}}_{k}^{2}-|{\tilde{u}}_{k}|^{2})^{2}+16\frac{p_{k}p_{2S+k}}{p_{k}+p_{2S+k}}{\tilde{q}}_{k}^{2}|{\tilde{u}}_{k}|^{2}\big{]}.$ (33)
It is interesting to draw our attention to the analogy in the structures, on the one hand, between Eqs. (IV.1) and (32) and, on the other hand, between Eqs. (22) and (33).
As with LQU, this makes it possible to actually calculate LQFI:
${\cal F}=\min{\{{\cal F}_{0},{\cal F}_{1}\}},$ (34)
where the branches (sub-functions) are equal to ${\cal F}_{0}=1-M_{zz}$ and ${\cal F}_{1}=1-M_{xx}$.
Substituting (11) into Eqs. (32) and (33), we come to a nice formula for the ${\cal F}_{0}$ branch
${\cal F}_{0}=4\sum_{k=1}^{2S}\frac{|u_{k}|^{2}}{a_{k}+a_{2S+k}}.$ (35)
Similarly, the final formula for the second branch ${\cal F}_{1}$ can be written as
${\cal F}_{1}=1-4\Bigg{\{}\frac{p_{0}(a_{2S+1}p_{0}+p_{1}p_{2S+1})}{(p_{0}+p_{1})(p_{0}+p_{2S+1})}+\frac{p_{4S+1}(a_{2S}p_{4S+1}+p_{2S}p_{4S})}{(p_{2S}+p_{4S+1})(p_{4S}+p_{4S+1})}+\frac{1}{2}\sum_{k=1}^{2S-1}\Bigg{[}\frac{(a_{k}p_{k+1}+p_{k}p_{2S+k})p_{k+1}}{(p_{k}+p_{k+1})(p_{k+1}+p_{2S+k})}$
$+\frac{a_{k}\{p_{2S+1+k}[p_{2S+k}(p_{k}+p_{k+1})+p_{k}p_{k+1}]+p_{k}p_{k+1}p_{2S+k}\}+p_{k}p_{2S+k}(p_{k}p_{2S+k}-p_{k+1}p_{2S+1+k})}{(p_{k}+p_{k+1})(p_{k+1}+p_{2S+k})(p_{2S+k}+p_{2S+1+k})(p_{2S+1+k}+p_{k})}(a_{2S+1+k}-a_{k+1})$
$+\frac{(a_{k}p_{2S+1+k}+p_{k}p_{2S+k})p_{2S+1+k}}{(p_{k}+p_{2S+1+k})(p_{2S+k}+p_{2S+1+k})}\Bigg{]}\Bigg{\}}.$ (36)
The resulting formulas (24), (25), (35) and (IV.2) for the branches of quantum correlations are surprisingly simple, compact and very comfortable to perform quick calculations on a computer.
V System at thermal equilibrium
Let’s now turn to the consideration of a spin system in the state of thermal equilibrium with a reservoir at the temperature $T$.
V.1 Spin Hamiltonian and Gibbs density matrix
We take the Hamiltonian in the Hermitian $\cal A$ form
${\cal H}=\left(\begin{array}[]{clccccccrc}E_{0}&&&&&&&&&\\ &h_{1}&0&\ldots&0&g_{1}&0&\ldots&0&\\ &0&h_{2}&\ldots&0&0&g_{2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&h_{2S}&0&0&\ldots&g_{2S}&\\ &g_{1}^{*}&0&\ldots&0&h_{2S+1}&0&\ldots&0&\\ &0&g_{2}^{*}&\ldots&0&0&h_{2S+2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&g_{2S}^{*}&0&0&\ldots&h_{4S}&\\ &&&&&&&&&E_{4S+1}\end{array}\right).$ (37)
The energy levels of the system are obviously $E_{0}$, $E_{4S+1}$ and $2S$ additional pairs
$E_{k,2S+k}=\frac{1}{2}\big{(}h_{k}+h_{2S+k}\pm R_{k}\big{)},$ (38)
where
$R_{k}=\sqrt{(h_{k}-h_{2S+k})^{2}+4|g_{k}|^{2}}$ (39)
and $k=1,\ldots,2S$.
The partition function $Z=\sum_{n}\exp(-E_{n}/T)$ of the system is given by
$Z=e^{-E_{0}/T}+e^{-E_{4S+1}/T}+2\sum_{k=1}^{2S}\cosh{\frac{R_{k}}{2T}}e^{-(h_{k}+h_{2S+k})/2T}.$ (40)
Statistical weights, i.e. density matrix eigenvalues (9), would be obtained as
$p_{0}=\frac{1}{Z}e^{-E_{0}/T},\qquad p_{k}=\frac{1}{Z}e^{-E_{k}/T},\qquad p_{2S+k}=\frac{1}{Z}e^{-E_{2S+k}/T},\qquad p_{4S+1}=\frac{1}{Z}e^{-E_{4S+1}/T}.$ (41)
The Gibbs density matrix is equal to
$\rho=\frac{1}{Z}\exp(-{\cal H}/T).$ (42)
Since $\cal H$ has the $\cal A$ form, this functional relationship also leads to the density matrix of the same structure, Eq. (3). If the dimension of the matrix $\cal H$ is finite, $\rho$ can be found manually or obtained on a digital machine by direct symbolic (analytical) calculations. After this, the results can be extended by induction to systems with arbitrary spin $S$. This allows us to get all non-zero matrix elements of the inner block of the density matrix (3) through the parameters of Hamiltonian:
$a_{k,2S+k}$ $=$ $\frac{1}{Z}\Big{(}\cosh{\frac{R_{k}}{2T}}\pm\frac{h_{2S+k}-h_{k}}{R_{k}}\sinh{\frac{R_{k}}{2T}}\Big{)}e^{-(h_{k}+h_{2S+k})/2T},$ (43)
$u_{k}=-\frac{2g_{k}}{ZR_{k}}\sinh{\frac{R_{k}}{2T}}e^{-(h_{k}+h_{2S+k})/2T}$
with $k=1,\ldots,2S$. In turn, these expressions open up an alternative opportunity to find the eigenvalues of the Gibbs density matrix using Eq. (9).
V.2 Hamiltonian and interaction constants
Expanding the Hermitian matrix (37) into generalized Gell-Mann matrices and then using the connections of these matrices with the spin ones, we obtain an expression for the Hamiltonian with $2(4S+1)$ terms of the spin matrices, their powers and products. Instead of this formal approach, one can also construct the same form of the Hamiltonian by adding linear independent spin terms with axial symmetry. In both cases, we come to the representation of the Hamiltonian (37) in the form
${\cal H}=B_{1}s_{z}+B_{2}S_{z}+J(s_{x}S_{x}+s_{y}S_{y})+J_{z}s_{z}S_{z}+D_{z}(s_{x}S_{y}-s_{y}S_{x})+K_{1}S_{z}^{2}+K_{2}s_{z}S_{z}^{2}+\ldots.$ (44)
We use a notation for the expanding coefficients which is suitable for the magnetic compounds, and will below restrict ourselves to seven parameters, namely, $B_{1}$ and $B_{2}$ are the $z$-components of an external magnetic field applied to spins 1/2 and $S$, respectively; $J$ and $J_{z}$ are the exchange Heisenberg constants; $K_{1}$ is the uniaxial one-ion anisotropy; $K_{2}$ is the uniaxial two-ion anisotropy; $D_{z}$ is the $z$-component of Dzyaloshinskii vector. It is interesting to note that the antisymmetric Dzyaloshinskii-Moriya (DM) interaction satisfies axial symmetry, while the symmetric combination $s_{x}S_{y}+s_{y}S_{x}$, which corresponds to the so-called Kaplan–Shekhtman–Entin-Wohlman–Aharony (KSEA) interaction, does not satisfy. Finally, if necessary, the Hamiltonian (44) can be extended by adding new axially invariant terms.
In Eq. (44), $s_{\alpha}=\sigma_{\alpha}/2$, where $\sigma_{\alpha}$ ($\alpha=x,y,z$) are again the Pauli matrices, and (setting $\hbar=1$)
$S_{z}={\rm diag}\,[S,S-1,\ldots,-S],$ (45)
$S_{x}=\left(\begin{array}[]{ccccccc}&b_{S}&&&&&\\ b_{S}&&b_{S-1}&&&&\\ &b_{S-1}&&b_{S-2}&&&\\ &&&&&&\\ &&&&&b_{-S+1}\\ &&&&b_{-S+1}&&\end{array}\right),$ (46)
$S_{y}=\left(\begin{array}[]{ccccccc}&-ib_{S}&&&&&\\ ib_{S}&&-ib_{S-1}&&&&\\ &ib_{S-1}&&-ib_{S-2}&&&\\ &&&&&&\\ &&&&&-ib_{-S+1}\\ &&&&ib_{-S+1}&&\end{array}\right)$ (47)
with $b_{m}=\sqrt{(S+m)(S-m+1)}/2$.
Inserting the Pauli and spin matrices (45) - (47) into Eq. (44) and then comparing the resulting open form of (44) with the matrix (37), we get expressions for the quantities $E_{0,4S+1}$, $h_{k,k+2S}$ and $g_{k}$ ($k=1,\dots,2S$) through the input parameters of the model, for each desired spin value $S$. For example, if $S=1$, then
$E_{0,5}=K_{1}+\frac{J_{z}+K_{2}}{2}\pm\Big{(}\frac{B_{1}}{2}+B_{2}\Big{)},\quad h_{1,4}=\pm\frac{B_{1}}{2},\quad h_{2,3}=-\frac{J_{z}}{2}+K_{1}\pm\Big{(}\frac{B_{1}+K_{2}}{2}-B_{2}\Big{)},\quad g_{1}=g_{2}=\frac{J+iD_{z}}{\sqrt{2}}.$ (48)
Now, one can get statistical weights (41) and entries (43), that opens up the possibility of calculating branches (24), (25), (35), (IV.2) and, finally, quantum correlations $\cal U$ and $\cal F$ through the interaction constants of the Hamiltonian (44).
VI Behavior of quantum correlations
Let us move on to consider the properties of spin-$(1/2,S)$ systems with a Hamiltonian (44). We will limit ourselves to seven specified interaction parameters because they are most important for magnetic materials.
VI.1 High-temperature wings of quantum correlations
Begin our analysis with the behavior of quantum correlations at high temperatures $T\to\infty$. Using symbolic calculations on a digital computer, we extract high-temperature asymptotics from the above general formulas. The results are as follows.
Spin $S=1$:
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{J^{2}+D_{z}^{2}}{3T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{3B_{1}^{2}+2K_{2}^{2}+2(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{12T^{2}}+O(1/T^{3}).$ (49)
Spin $S=3/2$:
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{5(J^{2}+D_{z}^{2})}{8T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{16B_{1}^{2}+41K_{2}^{2}+20(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{64T^{2}}+O(1/T^{3}).$ (50)
Spin $S=2$:
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{J^{2}+D_{z}^{2}}{T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{5B_{1}^{2}+34K_{2}^{2}+10(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{20T^{2}}+O(1/T^{3}).$ (51)
Spin $S=5/2$:
${\cal F}_{0}=2\,{\cal U}_{0}=\frac{35(J^{2}+D_{z}^{2})}{24T^{2}}+O(1/T^{3}),\quad{\cal F}_{1}=2\,{\cal U}_{1}=\frac{48B_{1}^{2}+707K_{2}^{2}+140(J^{2}+J_{z}^{2}+D_{z}^{2}+2B_{1}K_{2})}{192T^{2}}+O(1/T^{3}).$ (52)
These expressions show that the quantum correlations asymptotically fall to zero according to the power law $T^{-2}$. The field $B_{2}$ and strength of uniaxial single-ion anisotropy $K_{1}$ are absent here. The amplitudes of these high-temperature wings (coefficients in the $1/T^{2}$ law) demonstrate that the DM coupling $D_{z}$ and the Heisenberg interactions $J$ and $J_{z}$ enhance the quantum correlations, but the strength of these correlations decreases when the sign of the $B_{1}K_{2}$ product is negative. Note that the main high-$T$ terms ${\cal F}_{0,1}$ are twice as large as the corresponding terms ${\cal U}_{0,1}$.
VI.2 $S$ as a quantum correlation enhancer
($\rm\bf a$)Refer to caption($\rm\bf b$)
Figure 1: LQFI versus reduced temperature for the isotropic Heisenberg systems $(1/2,S)$ with $S=1$ (blue), $S=3/2$ (green), $S=2$ (orange) and $S=5/2$ (red). ($\rm\bf a$$J>0$, antiferromagnetic interactions; ($\rm\bf b$$J<0$, ferromagnetic interactions. Here, ${\cal H}=J{\vec{s}}\cdot{\vec{S}}$.
($\rm\bf a$)Refer to caption($\rm\bf b$)
Figure 2: LQFI versus reduced temperature for the isotropic Heisenberg systems $(1/2,S)$ with $S=1$ (blue), $S=3/2$ (green), $S=2$ (orange) and $S=5/2$ (red). ($\rm\bf a$$J>0$, antiferromagnetic interactions; ($\rm\bf b$$J<0$, ferromagnetic interactions. Here, ${\cal H}=\frac{J}{S}{\vec{s}}\cdot{\vec{S}}$.
Table 1: Values of ground-state quantum correlations LQU and LQFI in the spin-$(1/2,S)$ XXX dimers with both antiferromagnetic ($J>0$) and ferromagnetic ($J<0$) couplings.
$S$ $J>0$ $J<0$
1/2 1 1/3=0.333…
1 8/9=0.888… 4/9=0.444…
3/2 5/6=0.833… 1/2=0.5
2 4/5=0.8 8/15=0.533…
5/2 7/9=0.777… 5/9=0.555…
Now let us explore the behavior of quantum correlations throughout temperature region. We start with the simplest case of the completely isotropic Heisenberg XXX model setting $J_{z}=J$ and $B_{1}=B_{2}=K_{1}=K_{2}=D_{z}=0$, i.e.
${\cal H}=J{\bf s}\cdot{\bf S}.$ (53)
Calculations show that here ${\cal F}_{1}={\cal F}_{0}={\cal F}$ and ${\cal U}_{1}={\cal U}_{0}={\cal U}$. Using Eq. (35), we find LQFI for different values of spin $S$. So, when $S=1$, the quantum correlation $\cal F$ is equal to
${\cal F}^{(1)}(T)=\frac{16}{9}\frac{\sinh{(3J/4T)\tanh{(3J/4T)}}}{2+\exp{(3J/2T)}}e^{3J/4T}.$ (54)
In analogous manner, we obtain
${\cal F}^{(3/2)}(T)=\frac{5\sinh^{2}{(J/T)}}{4+4\cosh{(2J/T)}-\sinh{(2J/T)}},$ (55)
${\cal F}^{(2)}(T)=\frac{32\sinh^{2}{(5J/4T)}}{5[5+5\cosh{(5J/2T)}-\sinh{(5J/2T)}]},$ (56)
${\cal F}^{(5/2)}(T)=\frac{35}{9}\frac{\cosh{(3J/T)}-1}{6+6\cosh{(3J/T)}-\sinh{(3J/T)}}$ (57)
[…]
Figure 1 shows the behavior of the quantum correlation $\cal F$ as a function of the reduced temperature $T/J$ (or $T/|J|$) for both antiferromagnetic and ferromagnetic systems with $S=1,\dots,5/2$. Notice, ${\cal F}$ and ${\cal U}$ coincide at $T=0$. As seen from Fig. 1, the quantum correlations have finite values at zero temperature. With increasing the temperature, the correlations first maintain a quasi-stationary value and then monotonically tend to zero.
In the case of antiferromagnetic interaction, the ground-state quantum correlations decrease with increasing spin $S$ (see Fig. 1a and Table 1). A similar picture is observed for the entanglement [53, 73]. This was interpreted as follows [53]. At $S\to\infty$, the system becomes more and more classical and therefore the quantum correlations gradually weaken and completely disappear in the limit of infinitely large spin $S$.
This interpretation is intuitively attractive, but unfortunately, it fails in the case of ferromagnets. Really, as can be seen from Fig. 1b and Table 1 (column $J<0$), quantum correlations at zero temperature surprisingly increase as spin $S$ increases. This is due to the peculiarities of the ground state.
[…]
It is interesting to draw a parallel with the behavior of quantum entanglement in the XXX model (53). Using negativity [15] as a measure of entanglement, the authors [53] found that the threshold temperature is equal to
$T_{\rm th}=\frac{2S+1}{2\ln{(2S+2)}}.$ (58)
Above this temperature, quantum entanglement vanishes. Note that quantum entanglement is present only in the case of antiferromagnetic interactions ($J>0$) and is completely absent for ferromagnets ($J<0$). From (58), a striking result follows [53]
$\lim_{S\to\infty}T_{\rm th}=\infty.$ (59)
In other words, the threshold temperature may be arbitrarily high when $S$ is sufficiently large.
However, this conclusion is controversial. Indeed, before passing to the limit $S\to\infty$, it is necessary to renormalize the spin operators: $S_{\alpha}\to S_{\alpha}/\sqrt{S(S+1)}$ ($\alpha=x,y,z$) [74, 75]. Because of this renormalization, the Hamiltonian (53) takes the form
${\cal H}=\frac{J}{\sqrt{S(S+1)}}\,{\bf s}\cdot{\bf S}\simeq\frac{J}{S}\,{\bf s}\cdot{\bf S}$ (60)
and the threshold temperature (58) should be replaced by
$T_{\rm th}=\frac{2S+1}{2S\ln{(2S+2)}}.$ (61)
From here we come to another result, namely,
$\lim_{S\to\infty}T_{\rm th}=0.$ (62)
So, increasing the spin length $S$ reduces the threshold temperature and turns it to zero in the limit of an infinitely large spin $S$.
The renormalized Hamiltonian (60) changes the behavior of quantum correlations. The new curves are shown in Fig. 2 for $S>1/2$, i.e. for relatively large spins. The main difference is occurred in the high-temperature region. Now an increase in the spin $S$ here leads to a decrease in quantum correlations. This is entirely consistent with intuitive expectations.
It is important to emphasize the following. The low-temperature limit of quantum correlations does not depend on the value of the interaction constant in the Hamiltonian (depends only on the sign of $J$). This is clearly visible from Figs. 1 and 2. Therefore, the values presented in Table 1 are universal.
VI.3 Sadden change phenomena
Quantum correlations LQU and LQFI are determined by Eqs. (23) and (34), respectively. In general, these functions are piecewise-defined. The minimization conditions included in their definitions can lead to an interesting phenomenon, namely, the behavior of quantum correlations can undergo sharp changes as the parameters of the system change smoothly. In the mathematical literature, such phenomena are called catastrophes [76, 77].
Consider the system, in which $B_{1}=K_{2}=D_{z}=0$. As follows from Eqs. (49) - (52), 0-wings are below 1-wings if $|J_{z}|>1$, and, conversely, above if $|J_{z}|<1$.
In the absence of a second external field ($B_{2}=0$), pairs of branches ${\cal U}_{0}(T)$ and ${\cal U}_{1}(T)$, as well as ${\cal F}_{0}(T)$ and ${\cal F}_{1}(T)$ do not intersect. For $|J_{z}|>1$, the quantum correlations ${\cal U}(T)={\cal U}_{0}(T)$ and ${\cal F}(T)={\cal F}_{0}(T)$. If $|J_{z}|<1$, then the correlations ${\cal U}(T)$ and ${\cal F}(T)$ are completely determined by the branches ${\cal U}_{1}(T)$ and ${\cal F}_{1}(T)$, respectively. In both cases, the quantum correlations are smooth functions.
[…]
Figure 4: Quantum correlations $\cal F$ and $\cal U$ as functions of temperature $T$ for the model with $B_{1}=0$, $B_{2}=-0.8$, $J=1$, $J_{z}=0.3$, $K_{1}=0$, $K_{2}=0$, and $D_{z}=0$ at different values of spin length $S$.
The situation begins to change dramatically at $B_{2}\neq 0$. Figure 3 shows the temperature dependencies of branches for different spin values $S$. The intersections of branches are clearly observed here. The ${\cal U}_{0}(T)$ curves intersects the ${\cal U}_{1}(T)$ ones at the points with temperatures 0.215, 0.262, 0.292 and 0.323 for $S=1$, 3/2, 2 and 5/2, respectively. These temperatures increase with increasing $S$. Similarly for the $\cal F$-branches.
On the correlation curves, the above intersection points are visible as fractures (sharp bends), see Fig. 4. They are points of non-analyticity. The functions ${\cal U}(T)$ and ${\cal F}(T)$ are continuous, but their first derivatives experience discontinuities of the first kind and therefore these functions belong to the differentiability class $C^{0}$.
Experimental detection of singular points and study of the behavior of quantum correlations near them carry valuable information about the systems under examination. This information can make it possible to draw conclusions about interaction constants and understand the role of certain bonds in processes occurring in various substances.
Figure 5: Branches of quantum correlations depending on temperature for the model with $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K_{1}=0.4$, $K_{2}=-1.02$ and $D_{z}=-0.3$ at spin lengths $S=1$ and 3/2.
As follows from Figs. 3 and 4, the above systems contain one sudden transition. It is remarkable that a series of such transitions is possible. Indeed, let $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K_{1}=0.4$, $K_{2}=-1.02$ and $D_{z}=-0.3$. As can be seen from Fig. 5, the $\cal U$-branches (blue curves) intersect twice and hence the behavior of LQU suddenly changes at two different temperatures. It is interesting to note that the second correlation, LQFI, preserves continuity and smoothness. This indicates a peculiar hybrid behavior of the non-classical correlation in the given system.
Figure 6: Branches of quantum correlations depending on temperature for the model with $B_{1}=-0.65$, $B_{2}=0.9$, $J=-1.3$, $J_{z}=-1$, $K_{1}=0.4$, $K_{2}=-0.6$ and $D_{z}=-0.3$ at spin lengths $S=1$ and 3/2. Insets show the vicinities of intersection points at higher temperatures.
Finally, let’s consider a model with slightly different parameters, namely, $B_{1}$, $B_{2}$, $J$, $J_{z}$, $K_{1}$ and $D_{z}$ remain the same, but $K_{2}$ is replaced by -0.6. Look at Fig. 6. Here, each pair of 0- and 1-branches intersects at two points. As a result, when the temperature changes from zero to infinity, the quantum correlations $\cal U$ and $\cal F$ undergo two sudden transitions. Figure 6 clearly shows the intersections of the curves at low temperatures. The $\cal U$ branches (blue curves) intersect at temperatures 0.164 and 0.06 for $S$=1 and 3/2, respectively. Similarly for $\cal F$-branches (orange curves): they intersect at temperatures 0.375 and 0.13 for spin lengths $S$=1 and 3/2, respectively. Unfortunately, at higher temperatures the intersections become less visible, so their surroundings are shown in the insets. From these insets it is clear that at high temperatures the behavior of quantum correlations is determined by 0-branches. The same conclusion follows from the high-temperature decompositions (49) - (50).
VII Summary and outlook
In this paper, we have derived analytical formulas (24), (25), (35) and (IV.2) for the branches of quantum correlations LQU and LQFI. The obtained formulas are surprisingly compact and allow one to calculate the local quantum uncertainty and local quantum Fisher information for any spin-(1/2,$S$) systems whose states commute with the $z$-component of the total spin.
The derived formulas were then applied to the systems in thermal equilibrium, for which the behavior of quantum correlations has been studied in detail. Among findings, we note the following. It is natural to expect that an increase in the spin length, $S\to\infty$, should destroy quantumness of the system. Our calculations confirm this observation, which is known for antiferromagnets at ground states. However, contrary to intuitive expectations, we found enhancement of quantum correlations at zero temperature for systems with ferromagnetic couplings.
We also carefully studied the phenomena of sudden transitions where smooth variations in temperature lead to sharp changes in the behavior of quantum correlations. Remarkably, both single such transitions and series of them have been revealed.
The obtained formulas can be used to study other properties of quantum correlations, such as their dynamical behavior.
Acknowledgments: S. H. was supported by Semnan University under Contract No. 21270. M.A.Y. was supported in part by a state task, the state registration number of the Russian Federation is ##\##124013000760-0.
CRediT authorship contribution statement
Saeed Haddadi: Methodology, Investigation. M. A. Yurischev: Writing - original draft, Conceptualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.
ORCID iDs